when given 2 coordinates ie., (-3,-3) and (2,6) how do you know which one is y1,y2,x1,x2? which one would be put into the equation first as far as the top and bottom?
Answer by checkley77(12844)

(-3,-3) and (2,6)
You can assign them either way:
(x1,y1) and (x2,y2)
[Slope=(y2-y1)/(x2-x1)]
(6+3)/(2+3)=9/5
or:
(x2,y2) and (x1,y1)
(-3-6)/(-3-2)=-9/-5=9/5
Same answer.